When you begin with a neutral molecule, what charge results when you draw any relevant resonance structure?
Answer #1

When we start with neutral molecule, the resultant net charge in the relevant resonance structure is always equals to zero.

CH2=CH-CH=CH2    <----------> CH2--CH=CH-CH2+

             (I)                                                         (II)

In the above two resonating structures of 1,3-butadiene, the first structure neutral resonating structure and the net charge is zero whereas the structure II is the charge species and the resultant net charge is (+1)+(-1)=0.